With a certain tail wind a jet aircraft arrives at its destination, 1,890 miles away, in 3 hours. Flying against the same wind, the plane makes the return trip in 3 3/8 hours. Find the wind speed and the plane's airspeed.
Answer by checkley71(8403)

1,890/3=630 SPEED OF THE PLANE WITH THE WIND.
1,890/3.375=560 SPEED OF THE PLANE AGAINST THE WIND.
(630-560)/2=70/2=35 MPH FOR THE WIND.
630-35=595 SPEED OF THE PLANE
560+35=595 DITTO.